Ex 4.1 Class 7 Maths Question 3.
Solve the following equations by trial and error method:
(i) 5p + 2 = 17
(ii) 3m – 14 = 4

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
(i) 5p + 2 = 17 For p = 1, LHS = 5 × 1 + 2 = 5 + 2 = 7 ≠ 17 (RHS) For p = 2, LHS = 5 × 2 + 2 = 10 + 2 = 12 ≠ 17 (RHS) For p = 3, LHS = 5 × 3 + 2 = 15 + 2 = 17 = 17 (RHS) Since the given equation is satisfied for p = 3 Thus, p = 3 is the required solution. (ii) 3m – 14 = 4 For m = 1, LHS = 3 × 1 – 14 = 3 – 14 = -11 ≠ 4 (RHS) For m = 2, LHS = 3 × 2 – 14 = 6 – 14 = -8 ≠ 4 (RHS) For m = 3, LHS = 3 × 3 – 14 = 9 – 14 = -5 ≠ 4 (RHS) Form m = 4, LHS = 3 × 4 – 14 = 12 – 14 = -2 ≠ 4 (RHS) For m = 5, LHS = 3 × 5 – 14 = 15 – 14 = -1 ≠ 4 (RHS) For m = 6, LHS = 3 × 6 – 14 = 18 – 14 = 4 (=) 4 (RHS) . Since, the given equation is satisfied for m = 6. Thus, m = 6 is the required solution.